OpenStudy (anonymous):

lim(x->-2) (x+1)/((x^2)+4x+3)

OpenStudy (anonymous):

the denominator can be factored into (x+1)(x+3) so you can re-write it as lim (x-->-2) [(x+1)/(x+1)(x+3)] We can cancel the (x+1) binomials to get 1/(x+3) Now just substitute x for -2 and you should get lim (x-->-2) [1/(x+3)]=1/(-2+3)=1/1=1 Hope this helps :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

No problem ;)